The ages of two brothers are in the ratio two to three, but in eight years the ratio of their ages will be three to four. Find their present age?
Answer by stanbon(75887)

The ages of two brothers are in the ratio two to three, but in eight years the ratio of their ages will be three to four. Find their present age?
------------------
NOW DATA:
Let the two brothers ages be 2x and 3x.
-------------------
Eight yrs. from now DATA:
2x will be 2x+8
3x will be 3x+8
--------------------
Eight yrs from now EQUATION:
2x+8/3x+8 = 3/4
-------
Cross-multiply to get:
3(3x+8) = 4(2x+8)
9x+24 = 8x+32
x = 8
----------
2x = 16 (one of the brother's age)
3x = 24 (the other brother's age)
----------------------------------
Cheers,
Stan H.